# Barrier-Riccati Synthesis for Nonlinear Safe Control with Expanded Region of Attraction

### v2

@@ Abstract
We present a Riccati-based framework for safety-critical nonlinear control that integrates the barrier states (BaS) methodology with the State-Dependent Riccati Equation (SDRE) approach. The BaS formulation embeds safety constraints into the system dynamics via auxiliary states, enabling safety to be treated as a control objective. To overcome the limited region of attraction in linear BaS controllers, we extend the framework to nonlinear systems using SDRE synthesis applied to the barrier-augmented dynamics and derive a matrix inequality condition that certifies forward invariance of a large region of attraction and guarantees asymptotic safe stabilization. The resulting controller is computed online via pointwise Riccati solutions. We validate the method on an unstable constrained system and cluttered quadrotor navigation tasks, demonstrating improved constraint handling, scalability, and robustness near safety boundaries. This framework offers a principled and computationally tractable solution for synthesizing nonlinear safe feedback in safety-critical environments.
@@ I Introduction
Safe autonomous control is increasingly demanded in modern simple and complex control systems. Since constraints are ubiquitous in control systems, different techniques have been proposed in the literature to handle them. Early in the control literature, many of those techniques were based on rendering set invariance through Lyapunov analysis [blanchini1999set]. These methods were primarily developed for linear systems or for systems with simple constraints, such as box constraints, and did not generalize easily to nonlinear settings. However, with advancements in control theory and the aid of increased computational power, many of these approaches have been extended and further investigated for nonlinear systems. Barrier methods including barrier certificates [prajna2003barrier, prajna2004safety], control barrier functions (CBFs) [wieland2007constructive, ames2014control, romdlony2014uniting], Barrier-Lyapunov functions (BLFs) [tee2009barrier_Lyap_automatica], and most recently barrier states (BaS) [almubarak2021safeddp, almubarak2023BaSTheorey], have shown strong success in the recent safe control literature.
@@ I-A Background and Related Work
The barrier states (BaS) augmentation method enforces forward invariance of the set of allowed states through augmenting the state of safety by means of barrier functions into the model of the safety-critical control system. This creates a transformed problem with the objective of seeking a stabilizing control law for the augmented model [almubarak2023BaSTheorey]. Linear control methods, such as pole-placement and the linear quadratic regulator (LQR), were used to safely stabilize the safety-critical system [Almubarak2021SafetyEC, aoun2023l1]. The same concept was extended to trajectory optimization in [almubarak2021safeddp], where BaS augmentation was shown to consistently outperform penalty methods and CBF safety filters in providing safe optimal trajectories. However, a byproduct of this method is the increase in the nonlinearity of the system, including the conversion of linear models to nonlinear ones [almubarak2023BaSTheorey]. Hence, the validity of the controller, portrayed through the region of attraction of the closed-loop system, is limited. This calls for a practical nonlinear control that provides a sufficiently large region of attraction.
State-Dependent Riccati Equation (SDRE) approaches are practical and widely used methods for obtaining sub-optimal control in nonlinear dynamical systems. They are used across many industries for regulation, control, observation, estimation, and filtering of nonlinear systems [cimen2008state, nekoo2019SDREReview]. By factorizing the dynamics into state-dependent coefficient matrices, the nonlinear system is expressed in a form analogous to the standard linear state-space equation. A further advantage is that the formulation resembles a linear one, while the inherent nonlinearities of the original system remain present in the closed-loop dynamics. It is worth mentioning that the solution to the SDRE is not guaranteed to be optimal [nekoo2019SDREReview]. Local optimality, however, is guaranteed under the standard assumptions that the parameterization is pointwise stabilizable, detectable, and continuously differentiable [cloutier1996nonlinear].
Early efforts to integrate safety into SDRE-based control include introducing a fictitious output representing the rate of constraint violation, as in [cloutier2001state]. The idea was to penalize this output in the cost, steering it to zero to maintain constraint satisfaction. However, this approach requires the fictitious output dynamics to be invertible, limiting its applicability to only systems equal number of constraints and control inputs. A later method [devi2021barrier], inspired by Barrier Lyapunov Functions (BLF), penalizes constraint violations through a BLF in the state-dependent weights in the cost. Yet, it is restricted to box constraints and introduces non-quadratic, non-differentiable terms, with a flawed stability proof based on an incorrect assumption about the time-derivative of the Riccati solution.
@@ I-B Contributions and Organization
The main contributions of this paper are threefold. First, using the Mean Value Theorem with proper factorization, we derive the barrier state equation into an extended linearization form, resulting in safety embedded models that preserve the structure of the linear-like form needed in the SDRE technique. Second, we propose the safety-embedded State-Dependent Riccati Equation control framework as a nearly optimal, nonlinear control method with embedded safety guarantees. Third, a major contribution is the derivation of a matrix inequality condition that is shown to provide semi-global asymptotic safe stability if satisfied.
The paper is organized as follows. Section II presents the problem formulation and assumptions. Section III introduces barrier states, derives the extended BaS dynamics, and formulates the safety-embedded SDRE for approximating the constrained HJB solution, with theoretical guarantees for safe, near-optimal control. Conditions for extending local to semi-global safe stability are also given. Finally, simulation results are presented in Section IV, followed by concluding remarks in Section V.
@@ II-A Safety-Critical Optimal Control Problem
where $t\in\mathbb{R}$ is time [^1] For notational convenience, dependence on time $t$ is dropped throughout the paper., $x\in\mathcal{X}\subset\mathbb{R}^{n}$ is the state, $u\in\mathcal{U}\subset\mathbb{R}^{m}$ is the control input, $x_{0}=x(0)$ is the initial state, $\mathbf{Q}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{+}\ \forall x\neq 0$ is a state cost function, $R:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m\times m}\succ 0\ \forall x\neq 0$ is a control cost function, $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ and $g:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times m}$ are at least $C^{1}$ at which we assume that for the unforced system $f(0)=0$, without loss of generality, and that the linearized system is stabilizable and detectable.
@@
It is assumed that the system is safely stabilizable ([almubarak2023BaSTheorey, Definition 4]) in a set containing the origin, i.e., there exists a continuous controller $u(x)$, for which the origin of the closed-loop system $\dot{x}=f(x)+g(x)u(x)$ is asymptotically stable and that the safe set is controlled invariant.
@@ III Safety Embedded State Dependent Riccati Equation Control
Consider the optimal control problem (1)-(2). In the linear–quadratic regulator (LQR) setting, the optimal value function is quadratic and obtained via the algebraic Riccati equation (ARE), yielding a linear state feedback law. The state-dependent Riccati equation (SDRE) method extends this idea to nonlinear systems by factorizing the dynamics into a pseudo-linear form with state-dependent matrices, then solving a Riccati equation pointwise. This enables locally Riccati-based, near-optimal feedback while preserving the structure and intuition of the linear theory.
More specifically, using the idea of extended linearization, also known as state-dependent coefficient (SDC) parameterization, we have
@@
subject to (5), where $Q:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times n}\succeq 0$ and $R:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m\times m}\succ 0$ are at least $C^{1}$ and it is assumed that the pairs $\big(A(x),g(x)\big)$ and $\big(A(x),Q^{\frac{1}{2}}(x)\big)$ are point-wise stabilizable and detectable, respectively, $\forall x\in\Omega$ such that the solution to this problem, known as the value function $V(x)$, exists and is continuously differentiable. For a detailed treatment of the assumptions required for existence, uniqueness, stability, and near-optimality, see [cimen2008state].
III-A Factorized Barrier States and Extended Safety Embedded Systems
@@
Consequently, by definition, $\beta(x)\to\infty$ if and only if $h(x)\to 0$ (approaching unsafe regions) [almubarak2023BaSTheorey]. The barrier’s time derivative is then given by
@@
where $\gamma\in\mathbb{R}^{+}$, $\beta^{\circ}=\beta(0)$ and the shift by $\beta^{\circ}$ ensures that the BaS dynamics vanish at $(x=0,u=0)$. It is worth noting that the term scaled by $\gamma$ ensures stabilizability of the new system after the BaS is augmented [almubarak2023BaSTheorey]. Furthermore, it can easily be verified that the barrier state dynamics (8) are smooth, given the smoothness of the system dynamics, $\beta(x)$, and the barrier function $\mathbf{B}$. For detailed derivations and theoretical discussions on barrier states and related stability results, see [almubarak2023BaSTheorey].
By construction, the barrier state $z$ remains bounded if and only if the barrier function $\beta(x)$ remains bounded [almubarak2023BaSTheorey, Lemma 1]; therefore, ensuring boundedness of $z$ directly guarantees system safety. The barrier state augmentation thus transforms the original system into an extended nonlinear system, even if the original system dynamics and constraints were linear. Consequently, a practical nonlinear control approach, such as the State-Dependent Riccati Equation (SDRE) method, is particularly suitable. In what follows, we present an extended linearization formulation of this safety-embedded augmented system that fits naturally within the SDRE framework.
@@
where $\tilde{\beta}(x)=(\int_{0}^{1}\frac{\partial\beta}{\partial x}(\mu x)d\mu)x$ by the Mean Value Theorem. Using this factorization, the barrier state dynamics (8) can equivalently be expressed in an extended linearization form as
$\dot{z}$ $=\underbrace{\left(\mathbf{B}^{\prime}\big(\mathbf{B}^{-1}(z+\beta^{\circ})\big)L_{A(x)}h(x)+\gamma\tilde{\beta}(x)\right)}_{A^{z}(x,z)}x$ (10)
@@
We refer to (11) as the extended safety embedded dynamics. One can see how similar the structure of the extended state-dependent matrices (12) is to the linear system’s ones presented in [almubarak2023BaSTheorey]. Here, we present a single BaS for simplicity. In general, for a system with multiple ($\mathcal{Q}$) constraints, one can either design a single aggregated barrier state using an aggregated safety function defined by
$\frac{1}{h}=\sum_{i=1}^{\mathcal{Q}}\frac{1}{h_{i}},$
or introduce multiple barrier states separately, as shown in [almubarak2021safeddp] or [almubarak2023BaSTheorey], depending on the problem and a desired performance. It is worth noting that this system has its origin as an equilibrium point and $\bar{f}(0)=0$ preserves the smoothness and stabilizability of the original control system (2) [Almubarak2021SafetyEC]. Therefore, the safety constraint is embedded in the closed-loop system’s dynamics, and stabilizing the safety embedded system (11) implies enforcing safety for the safety-critical system (2), i.e., forward invariance of the safe set $\mathcal{S}$ with respect to (2). This crucial result is formally established in [almubarak2023BaSTheorey, Theorem 1], and a direct implication of the theorem was discussed in [almubarak2023BaSTheorey, Remark 2]. This result is depicted in the following corollary.
Corollary 1 ([almubarak2023BaSTheorey, Theorem 1]).
Let $u=K(\bar{x})$ be a continuous feedback controller such that the origin of the safety embedded closed-loop system, $\dot{\bar{x}}=\bar{A}(\bar{x})\bar{x}+\bar{g}(\bar{x})K(\bar{x})$, is asymptotically stable. Then, there exists an open set $\mathcal{A}_{\text{safe}}\subseteq\mathcal{S}$ such that $u=K(\bar{x})$ is safely stabilizing $\forall x(0)\in\mathcal{A}_{\text{safe}}\subseteq\mathcal{S}$.
@@ III-B Safety Embedded SDRE Control
subject to (11), where $Q:\mathbb{R}^{n+q}\rightarrow\mathbb{R}^{{n+q}\times{n+q}}\succeq 0$, $R:\mathbb{R}^{n+q}\rightarrow\mathbb{R}^{m\times m}\succ 0\ \forall\bar{x}\neq 0$, and $q$ is the number of constructed barrier states. Note that $Q$ needs to be chosen such that the pair $(\bar{A},Q^{\frac{1}{2}})$ is detectable after adding the BaS [cimen2008state]. It is worth noting that with this formulation, the $Q$ matrix can be a simple real-valued constant matrix which ensures the convexity of the cost function. That is, the BaS can be penalized as any other state and hence the proposed method provides a natural way to incorporate constraints. By Corollary 1, solving this optimal control problem implies solving both the safety and stability problems.
@@
$\frac{\partial V}{\partial\bar{x}}^{\top}=P(\bar{x})\bar{x},\ P:\mathbb{R}^{n+q}\rightarrow\mathbb{R}^{{n+q}\times{n+q}}\succ 0$ (15)
where $P(\bar{x})$ solves the safety embedded state-dependent Riccati equation, that we refer to as BaS-SDRE, which is given by

### v3

@@ Abstract
We present a Riccati-based framework for safety-critical nonlinear control that integrates the barrier states (BaS) methodology with the State-Dependent Riccati Equation (SDRE) approach. The BaS formulation embeds safety constraints into the system dynamics via auxiliary states, enabling safety to be treated as a control objective. To overcome the limited region of attraction in linear BaS controllers, we extend the framework to nonlinear systems using SDRE synthesis applied to the barrier-augmented dynamics and derive a matrix inequality condition that certifies forward invariance of a large region of attraction and guarantees asymptotic safe stabilization. The resulting controller is computed online via point-wise Riccati solutions. We validate the method on an unstable constrained system and cluttered quadrotor navigation tasks, demonstrating improved constraint handling, scalability, and robustness near safety boundaries. This framework offers a principled and computationally tractable solution for synthesizing nonlinear safe feedback in safety-critical environments.
@@ I Introduction
Safe autonomous control is increasingly demanded in modern simple and complex control systems. Since constraints are ubiquitous in control systems, different techniques have been proposed in the literature to handle them. Early in the control literature, many of those techniques were based on rendering set invariance through Lyapunov analysis [7]. These methods were primarily developed for linear systems or for systems with simple constraints, such as box constraints, and did not generalize easily to nonlinear settings. However, with advancements in control theory and the aid of increased computational power, many of these approaches have been extended and further investigated for nonlinear systems. Barrier methods including barrier certificates [15, 14], control barrier functions (CBFs) [20, 5, 17], Barrier-Lyapunov functions (BLFs) [18], and most recently barrier states (BaS) [4, 3], have shown strong success in the recent safe control literature.
@@ I-A Background and Related Work
The barrier states (BaS) augmentation method enforces forward invariance of the set of allowed states through augmenting the state of safety by means of barrier functions into the model of the safety-critical control system. This creates a transformed problem with the objective of seeking a stabilizing control law for the augmented model [3]. Linear control methods, such as pole-placement and the linear quadratic regulator (LQR), were used to safely stabilize the safety-critical system [2]. The same concept was extended to trajectory optimization in [4], where BaS augmentation was shown to consistently outperform penalty methods and CBF safety filters in providing safe optimal trajectories. Additionally, extension of BaS to adaptive control methods was studied in [6, 1]. However, a byproduct of this method is the increase in the nonlinearity of the system, including the conversion of linear models to nonlinear ones [3]. Hence, the validity of the controller, portrayed through the region of attraction of the closed-loop system, is limited. This calls for a practical nonlinear control that provides a sufficiently large region of attraction.
State-Dependent Riccati Equation (SDRE) approaches are practical and widely used methods for obtaining sub-optimal control in nonlinear dynamical systems. They are used across many industries for regulation, control, observation, estimation, and filtering of nonlinear systems [8, 16]. By factorizing the dynamics into state-dependent coefficient matrices, the nonlinear system is expressed in a form analogous to the standard linear state-space equation. A further advantage is that the formulation resembles a linear one, while the inherent nonlinearities of the original system remain present in the closed-loop dynamics. It is worth mentioning that the solution to the SDRE is not guaranteed to be optimal [16]. Local optimality, however, is guaranteed under the standard assumptions that the parameterization is point-wise stabilizable, detectable, and continuously differentiable [11].
Early efforts to integrate safety into SDRE-based control include introducing a fictitious output representing the rate of constraint violation, as in [10]. The idea was to penalize this output in the cost, steering it to zero to maintain constraint satisfaction. However, this approach requires the fictitious output dynamics to be invertible, limiting its applicability to systems with an equal number of constraints and control inputs. A later method [12], inspired by BLFs, penalizes constraint violations through a BLF in the state-dependent weights in the cost. Yet, it is restricted to box constraints and introduces non-quadratic, non-differentiable terms, with a flawed stability proof based on an incorrect assumption about the time-derivative of the Riccati solution.
@@ I-B Contributions and Organization
The main contributions of this paper are threefold:
  • First, using the Mean Value Theorem with proper factorization, we derive the barrier state equation into an extended linearization form, resulting in safety embedded models that preserve the structure of the linear-like form needed in the SDRE technique.
  • Second, we propose the safety-embedded State-Dependent Riccati Equation control framework as a nearly optimal, nonlinear control method with embedded safety guarantees.
  • Third, a major contribution is the derivation of a matrix inequality condition that is shown to provide semi-global asymptotic safe stability if satisfied.
The paper is organized as follows. Section II presents the problem formulation. Section III introduces BaS, derives the extended BaS dynamics, and formulates the safety-embedded SDRE for approximating the constrained HJB solution, with theoretical guarantees for safe, near-optimal control. Conditions for extending local to semi-global safe stability are also given. Finally, simulation results are presented in Section IV, followed by conclusions in Section V.
@@ II-A Safety-Critical Optimal Control Problem
where $t\in\mathbb{R}$ is time [^1] For notational convenience, dependence on time $t$ is dropped throughout the paper., $x\in\mathcal{X}\subset\mathbb{R}^{n}$ is the state, $u\in\mathcal{U}\subset\mathbb{R}^{m}$ is the control input, $x_{0}=x(0)$ is the initial state, $\mathbf{Q}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{+}\ \forall x\neq 0$ is a state cost function, $R:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m\times m}\succ 0\ \forall x\neq 0$ is a control cost function, $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ and $g:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times m}$ are at least $C^{1}$, and we assume that for the unforced system $f(0)=0$, without loss of generality, and that the linearized system is stabilizable and detectable.
@@
It is assumed that the system is safely stabilizable ([3, Definition 4]) in a set containing the origin, i.e., there exists a continuous controller $u(x)$, for which the origin of the closed-loop system $\dot{x}=f(x)+g(x)u(x)$ is asymptotically stable and that the safe set is controlled invariant.
@@ III Safety Embedded State Dependent Riccati Equation Control
Consider the optimal control problem (1)-(2). In the linear–quadratic regulator (LQR) setting, the optimal value function is quadratic and obtained via the algebraic Riccati equation (ARE), yielding a linear state feedback law. The state-dependent Riccati equation (SDRE) method extends this idea to nonlinear systems by factorizing the dynamics into a pseudo-linear form with state-dependent matrices, then solving a Riccati equation point-wise. This enables locally Riccati-based, near-optimal feedback while preserving the structure and intuition of the linear theory.
Specifically, using extended linearization, also known as state-dependent coefficient (SDC) parameterization, we have
@@
subject to (5), where $Q:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times n}\succeq 0$ and $R:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m\times m}\succ 0$ are at least $C^{1}$ and it is assumed that the pairs $\big(A(x),g(x)\big)$ and $\big(A(x),Q^{\frac{1}{2}}(x)\big)$ are point-wise stabilizable and detectable, respectively, $\forall x\in\Omega$ such that the solution to this problem, known as the value function $V(x)$, exists and is continuously differentiable. For a detailed treatment of the assumptions required for existence, uniqueness, stability, and near-optimality, see [8].
III-A Factorized BaS and Extended Safety Embedded Systems
@@
Consequently, by definition, $\beta(x)\to\infty$ if and only if $h(x)\to 0$ (approaching unsafe regions) [3]. The barrier’s time derivative is then given by
@@
where $\gamma\in\mathbb{R}^{+}$, $\beta^{\circ}=\beta(0)$ and the shift by $\beta^{\circ}$ ensures that the BaS dynamics vanish at $(x=0,u=0)$. It is worth noting that the term scaled by $\gamma$ ensures stabilizability of the new system after the BaS is augmented [3]. Furthermore, it can easily be verified that the BaS dynamics (8) are smooth, given the smoothness of the system dynamics, $\beta(x)$, and the barrier function $\mathbf{B}$. For detailed derivations and theoretical discussions on BaS and related stability results, see [3].
By construction, the barrier state $z$ remains bounded if and only if the barrier function $\beta(x)$ remains bounded [3, Lemma 1]; therefore, ensuring boundedness of $z$ directly guarantees system safety. The barrier state augmentation thus transforms the original system into an extended nonlinear system, even if the original system dynamics and constraints were linear. Consequently, a practical nonlinear control approach, such as the SDRE method, is particularly suitable. In what follows, we present an extended linearization formulation of this safety-embedded augmented system that fits naturally within the SDRE framework.
@@
where $\tilde{\beta}(x)=\int_{0}^{1}\frac{\partial\beta}{\partial x}(\mu x)d\mu$ by the Mean Value Theorem. Using this factorization, the barrier state dynamics (8) can equivalently be expressed in an extended linearization form as
$\dot{z}$ $=\underbrace{\left(\mathbf{B}^{\prime}\big(\mathbf{B}^{-1}(z+\beta^{\circ})\big)L_{A(x)x}h(x)+\gamma\tilde{\beta}(x)\right)}_{A^{z}(x,z)}x$ (10)
@@
We refer to (11) as the extended safety embedded dynamics. The structure of the extended state-dependent matrices (12) is similar to the linear systems presented in [3]. Here, we present a single BaS for simplicity. In general, for a system with multiple ($\mathcal{Q}$) constraints, one can either design a single aggregated barrier state using an aggregated safety function defined by $\frac{1}{h}=\sum_{i=1}^{\mathcal{Q}}\frac{1}{h_{i}}$, or introduce multiple BaS separately, as shown in [4] or [3], depending on the problem and a desired performance. Note that this system has its origin as an equilibrium point: $\bar{A}(\bar{x})\bar{x}=0$, when $\bar{x}=0$. This preserves the smoothness and stabilizability of the original control system (2) [2]. Therefore, the safety constraint is embedded in the closed-loop system’s dynamics, and stabilizing the safety embedded system (11) implies enforcing safety for the safety-critical system (2), i.e., forward invariance of the safe set $\mathcal{S}$ with respect to (2). This crucial result is formally established in [3, Theorem 1], and a direct implication of the theorem was discussed in [3, Remark 2]. This result is formalized in the following corollary.
Corollary 1 ([3, Theorem 1]).
Let $u=K(\bar{x})$ be a continuous feedback controller such that the origin of the safety embedded closed-loop system, $\dot{\bar{x}}=\bar{A}(\bar{x})\bar{x}+\bar{g}(\bar{x})K(\bar{x})$, is asymptotically stable. Then, there exists an open set $\mathcal{A}_{\text{safe}}\subseteq\mathcal{S}$ such that $u$ is safely stabilizing $\forall x(0)\in\mathcal{A}_{\text{safe}}\subseteq\mathcal{S}$.
@@ III-B Safety Embedded SDRE Control
subject to (11), where $Q:\mathbb{R}^{n+q}\rightarrow\mathbb{R}^{{(n+q)}\times{(n+q)}}\succeq 0$, $R:\mathbb{R}^{n+q}\rightarrow\mathbb{R}^{m\times m}\succ 0\ \forall\bar{x}\neq 0$, and $q$ is the number of constructed barrier states. Note that $Q$ needs to be chosen such that the pair $(\bar{A},Q^{\frac{1}{2}})$ is detectable after adding the BaS [8]. It is worth noting that with this formulation, the $Q$ matrix can be a simple real-valued constant matrix which ensures the convexity of the cost function. That is, the BaS can be penalized as any other state and hence the proposed method provides a natural way to incorporate constraints. By Corollary 1, solving this optimal control problem implies solving both the safety and stability problems.
@@
$\frac{\partial V}{\partial\bar{x}}^{\top}=P(\bar{x})\bar{x},\ P:\mathbb{R}^{n+q}\rightarrow\mathbb{R}^{{n+q}\times{n+q}}$ (15)
where $P(\bar{x})\succ 0$ solves the safety embedded state-dependent Riccati equation, that we refer to as BaS-SDRE, which is given by